Let vector "a = (1, sqrt(2) ,1)" and vector "b = (2,0,2)"

  1. Find the angle "\theta" between the 2 vectors.
  2. Find the projection of a in the direction of b
1
Expert's answer
2021-07-06T08:28:48-0400

"let\space a=(1,\sqrt2,1)\\b=(2,0,2)\\"

a.

"cos(\theta)=\frac{a.b}{|a|.|b|}\\"

"=\frac{(1,\sqrt2,1).(2,0,2)}{\sqrt{1+2+1}{\sqrt{4+4}}}"


"=\frac{2+2}{\sqrt4 \sqrt8}=\frac{4}{2.2\sqrt2}"


"=\frac{1}{\sqrt2}"


"\theta=cos^{-1}(\frac{1}{\sqrt2})\\\theta=45\degree"



b.

Projection of a in direction of b

"=\frac{\vec{a}.\vec{b}}{|\vec{b}|}\space or\space |\vec{a}|cos\theta"


"\&\space |\vec{b}|=\sqrt8"


"=\frac{4}{\sqrt8}=\frac{4}{2\sqrt2}=\frac{2}{\sqrt2}=\sqrt2"